OpenStudy (anonymous):

solve: log(x+4)=log(5x+1)+log2

OpenStudy (anonymous):

x=2/9

OpenStudy (anonymous):

I don't understand how to do the work. How did you get that answer?

OpenStudy (anonymous):

log(x+4)=log2(5x+1) x+4=10x+2 (take anti log both sides) ...

OpenStudy (anonymous):

Ok, thank you, now I understand.